# Discovery and Analysis of Afterglows from Poorly Localised GRBs with the Gravitational-wave Optical Transient Observer (GOTO) All-sky Survey

### v1

@@ 4 Afterglow Follow-up Observations
Figure 5(a–g) illustrates the timeline of the multi-wavelength follow-up campaigns for all GRBs in the sample, marking the epochs at which each facility recorded its first observation relative to the trigger time. The various phases of the events are colour-coded as follows: X-ray afterglow (blue), UV/optical/NIR afterglow (green), radio afterglow (purple), and prompt emission (red) for completeness. This timeline highlights the wide temporal coverage and the rapid, coordinated response from both ground- and space-based observatories across the electromagnetic spectrum for the GRBs in our sample. More observation details are provided below. A full log of photometric measurements, combining GCN Circular reports and data from this work, is presented in Tables LABEL:tab:phot_data and A2, while Table 3 summarises the spectroscopic campaigns.
@@ 4.2.1 Swift/UVOT
In addition to XRT, Swift simultaneously observes with its UV-Optical Telescope (UVOT; Roming et al. 2005). We obtained the resulting data from the UK Swift Science Data Centre [^6] https://www.swift.ac.uk/index.php (UKSSDC) and used uvotproduct v2.9777As part of HEASoft v6.32 (Nasa High Energy Astrophysics Science Archive Research Center (Heasarc), 2014). to measure the photometry of the afterglow. We used a 5″ radius circular aperture centred at the positions noted in Section 3.2 and a detection threshold of $3\sigma$. The measured magnitudes were converted from the UVOT photometric system to AB using the standard UVOT zeropoints (Breeveld et al., 2011). The afterglow was detected in at least one epoch for five of the seven sources in our sample. The exceptions are GRB 240112A, where the UVOT FoV didn’t cover the source position, and GRB 240225B, for which no UVOT data were obtained. The UVOT follow-up observations are listed in Table LABEL:tab:phot_data, and the corresponding light curves are presented in Figure 6.
@@
The IO:O Imager at the robotic 2m Liverpool Telescope (LT; Steele et al., 2004) located at the international Observatorio del Roque de los Muchachos, La Palma, was triggered for follow-up of GRBs 240122A, 240225B, and 240910A. For GRBs 240122A and 240910A, one epoch was obtained in each of rizriz bands. For GRB 240225B, 1 epoch is in grizgriz, and three later epochs were secured in $r$-band. All LT data were pre-reduced for bias, dark, and flat-field corrections using the facility pipeline. The photometry was then extracted with the photometry-sans-frustration pipeline (psf; Nicholl et al., 2023), making use of its built-in template subtraction. The observations are summarised in Table LABEL:tab:phot_data, and the corresponding light curves are shown in Figure 6.
@@ 4.2.3 NOT
4.2.4 VLT/Xshoter acquisition camera
@@ 4.2.6 OSN
We processed all our observations using the APEX pipelines (Devyatkin et al., 2010; Kouprianov, 2012; Pankov et al., 2022). The process involved image calibration (dark frame subtraction, flat-fielding, cosmic-ray removal), image quality control, image stacking, and source extraction. The apex_forced_phot pipeline was utilised for the forced photometry of the GRB 240225B afterglow on difference images. Image subtraction with the method described in Tomaney & Crotts 1996, was performed by apex_subtract pipeline, using PanSTARRS-DR1 survey images as a reference obtained from HIPS2FITS service [^11] https://alasky.cds.unistra.fr/hips-image-services/hips2fits (Boch et al., 2020). This step ensured that the underlying host galaxy, presented in the Legacy Survey DR9 with a magnitude of $r\sim 24.2$ and a photometric redshift of $z\sim 0.9$, did not affect the afterglow measurement.
@@ 4.2.11 Altay
The 1.6m Multi-channel Photometric Survey Telescope (Mephisto) is a wide-field multichannel telescope (Yuan et al., 2020). It is located at Lijiang Observatory of Yunnan Astronomical Observatories, Chinese Academy of Sciences, and is operated by the South-Western Institute for Astronomy Research, Yunnan University. Equipped with three-channel CCD cameras (blue $uv$, yellow $gr$, and red $iz$ channels), Mephisto can perform simultaneous observations in $ugi_{m}$ or $vrz_{m}$ optical bands at a particular moment. The wavelength coverage of the $u_{m},v_{m},g_{m},r_{m},i_{m}$, and $z_{m}$ filters is 320–365, 365–405, 480–580, 580–680, 775–900, and 900–1050 nm with central wavelengths at 345, 385, 529, 628, 835, and 944 nm, respectively (see, e.g., Yang et al., 2024; Cheng et al., 2025). Presently, the facility is in an advanced stage of commissioning.
@@ 4.3 Spectroscopic Observations
For the GRBs in our sample, we acquired spectra using the X-shooter instrument mounted on the VLT (Vernet et al., 2011), and OSIRIS (Optical System for Imaging and lowIntermediate-Resolution Integrated Spectroscopy) on the Gran Telescopio Canarias (GTC; Cepa 1998), except for GRB 241002B, for which we did not get any spectroscopic observation due to scheduling constraints. A complete summary of the VLT/X-shooter and GTC/OSIRIS spectroscopic configurations, exposure details, and observing conditions is provided in Table 3, whereas spectra are shown in Figures (A1)–(A7).
@@ 4.4.1 ATCA
Following its optical localisation, GRB 240122A was observed with the ATCA under the PanRadio GRB programme C3542 (PI: Anderson) on 2024 January 24, 26, 28, and February 12. This program aims to perform high-cadence multi-frequency radio monitoring of a large sample of LGRBs in the southern hemisphere (Declinations <10<-10 deg) between minutes to years after the burst to explore the evolution and properties of their afterglows (Leung et al., in prep.; Anderson et al., in prep.). GRB 240122A was observed with a wide range of frequencies centred on 5.5, 9.0, 16.7, and 21.2 GHz, each with a 2048 MHz-wide band. We reduced the visibility data using standard routines in MIRIAD (Sault et al., 1995). We used a combination of manual and automatic RFI flagging before calibration, conducted with MIRIAD tasks uvflag and pgflag, respectively. We used PKS 1934-63 to determine the bandpass response and to calibrate the flux density scale for all frequency bands. We used PKS 0607-157 to calibrate the time-variable complex gains for all epochs and frequency bands. After calibration, we inverted the visibilities using a robust weighting of 0.5 and then used the CLEAN algorithm (Clark, 1980) to the target source field using standard MIRIAD tasks INVERT, CLEAN, and RESTOR to obtain the final images. For each observation, we measure the flux density of a detected source by fitting a point-source model to the restored image using the Miriad task imfit and report a non-detection using the rms sensitivity obtained from the residual image. The $1\sigma$ errors reported are purely statistical, as the systematic errors are expected to be much smaller ($\lesssim 5$ per cent; e.g., Reynolds, 1994; Tingay et al., 2003). We detected the radio counterpart at 9 GHz on both 2024 January 24 and 28 at a position consistent with the GOTO optical counterpart (Kumar et al., 2024b). For all other frequencies, we estimated $3\sigma$ upper limits (see Table A2).
@@ 4.4.2 VLA
We observed GRB 240910A with the VLA 3.1 (2024 September 13, Giarratana et al. 2024a), 9.1 (2024 September 19), 21.3 (2024 October 1) and 46.2 (2024 October 26) days post-burst (Project code: SF171028) at the central frequencies of 6 (C band), 10 (X band) and 15 GHz (Ku band), with a bandwidth of 4, 4 and 6 GHz, respectively. The VLA source J0125-0005 was used as a phase calibrator. The distance between the target and the phase calibrator was about 2.7. Each observation included scans on the flux and bandpass calibrator 3C48. The data were calibrated using the custom casa pipeline (Version 6.5.4; McMullin et al., 2007) and visually inspected for possible radio frequency interference. The final images were produced with the tclean task in casa (Version 6.5.4) using a Briggs weighting scheme (robust =0.5=0.5). Results from the campaign are reported in Table A2. The GRB is detected at all frequencies during the first two epochs, and at 6 GHz it is also detected in the third epoch. The maximum flux densities, measured in the first epoch, are $37\pm 10, 11$$14\pm 9$, and $86\pm 10$ uJy at 6, 10, and 15 GHz, respectively. For each detection, the flux density was measured by fitting a Gaussian to the cleaned image using the imview task in casa. The final flux density error was estimated as the squared sum of the root mean square (RMS) and a typical 5% accuracy for the amplitude scale calibration. Upper limits are reported with a $5\sigma$ confidence level.
@@
GRB $z$ $T_{\rm 90}$ Model Fluence [2 – 20 keV] $\alpha$ $\beta$ $E_{\rm p}$ PG-Statistic DoF HR
@@
240122A $3.163$ $\approx 36$ PL $1.26^{+0.47}_{-0.94}$ $1.9\pm 0.4$ - - $1.32^{+1.30}_{-0.65}$
240225B $0.946$ $\approx 21$ PL $10.7^{+0.63}_{-1.26}$ $1.9\pm 0.1$ - - $1.32^{+0.25}_{-0.21}$
GRB $z$ $T_{\rm 90}$ Model Fluence [10 – 1000 keV] $\alpha$ $\beta$ $E_{\rm p}$ PG-Statistic DoF HR
@@ 5.1 Prompt Emission Analysis
Here, we describe the methods used to extract and analyse the prompt emission properties for each GRB, using data from the relevant high-energy instruments. The analysis is divided into two parts: the MAXI/GSC events and the Fermi/GBM events.
@@ 5.1.1 GRBs 240122A and 240225B with MAXI/GSC
We analysed the prompt emission of GRBs 240122A and 240225B with archival data of MAXI/GSC using High Energy Astrophysics Software (HEASOFT [^14] https://heasarc.gsfc.nasa.gov/docs/software/heasoft/). X-ray events of gsc_med type are processed with mxproduct. Because light curves and spectra produced by mxproduct are not suitable for short and variable transients like GRBs, we performed an additional process to extract light curves with a 1-second time resolution and applied effective area correction. Note that the process is identical to the process used for the MAXI GRB catalogue.
@@
Then, we extract a spectrum of the $T_{\rm 90}$ interval and correct for variations in the effective area. The spectra of GRBs 240122A and 240225B are fit with a single power-law model, and the photon indices are found to be $1.9\pm 0.4$ and $1.9\pm 0.1$ ($1\sigma$ error), respectively. The energy flux in 2–20 keV was $3.5_{-2.6}^{+1.3}\times$10-9 erg cm-2 s-1 and $5.1_{-0.6}^{+0.3}\times$10-8 erg cm-2 s-1. Figure A9 shows the spectrum with the best-fit model. The results of each fit are shown in Table 4.
@@ 5.1.2 GRBs 240619A, 240910A, 240916A, 241002B, and 241228B with Fermi/GBM
We analysed the prompt emission of GRBs 240619A, 240910A, 240916A, 241002B, and 241228B using the Fermi/GBM data available from the High Energy Astrophysics Science Archive Research Center (HEASARC [^15] https://heasarc.gsfc.nasa.gov) archive (von Kienlin et al., 2020; Gruber et al., 2014; von Kienlin et al., 2014; Bhat et al., 2016). Using the HEASOFT and the Fermi Gamma-ray Data Tools (Goldstein et al., 2023), we took the time-tagged event (TTE) data in the $T_{\rm 90}$ interval for each burst to use as the source. We used the data from the brightest NaI detectors and the corresponding BGO detectors. These were NaI 0, NaI 1, NaI 2 and BGO 0 for GRB 240619A; NaI 8, NaI 11, and BGO 1 for GRB 260910A; NaI 3, NaI 4, and BGO 0 for GRB 240916A; NaI 10, NaI 11, and BGO 1 for GRB 241002B; and NaI 6, NaI 7, and BGO 1 for GRB 241228B. The background was modelled in the standard way, using a polynomial function fit to the CSPEC data for each burst. The order of the polynomial was increased until the reduced fit statistic was <1.15<1.15, resulting in the models fitting each GRB suitably well. Once the background model was obtained, we interpolated it across the source interval and exported this as the background to be used for spectral analysis. Additionally, we extracted the necessary response files.
@@ 5.2 Prompt Emission Properties
We plot the 4/5 Fermi bursts where it was possible to measure both $E_{\rm p}$ and $E_{iso}$ (see Table 4) on the Amati plane (Amati et al., 2002; Amati, 2006) in Figure 7. GRBs 240916A and 241228B appear to show unusually high redshift-corrected peak energies ($E_{p,i}$) relative to their $E_{iso}$ measurements, and are inconsistent with the Amati relation at the $3\sigma$ level. Both bursts lie at the high end of the $E_{iso}$ distribution, indicating intrinsically powerful GRBs. In addition to this, the high energy photon index ($\beta$) measured for GRB 240619A is less than 2, indicating that the true peak of the spectrum is at an even higher energy. This may make 240619A an outlier of the Amati relation too. Only GRB 240910A appears to be typical in terms of its measured prompt properties.
We also plot all five Fermi and two MAXI GRBs on an HR – $T_{\rm 90}$ diagram (c.f. Kouveliotou et al., 1993), see Figure 8. It is immediately apparent that the Fermi GRBs are unusually spectrally hard. Only GRB 241002B (best fit with a power-law model with a photon index of $\alpha=1.97^{+0.08}_{-0.06}$) sits within the main ‘cloud’ of LGRBs. The others are either longer in $T_{\rm 90}$ (GRB 240910A), or driven to high HR by their abnormally hard (for collapsar GRBs) photon indices (GRBs 240619A, 241228B) and/or high $E_{\rm p}$ (GRB 240916A). The MAXI GRBs also sit at the high end of the GBM HR distribution, but these values are obtained by extrapolating the 2–20 keV spectral fits to 300 keV, and so should be considered upper limits because we have no constraints on any breaks in the spectrum.
Figure 7: Fermi GRBs in our sample plotted in the Amati plane (Amati et al., 2002; Amati, 2006), showing the relationship between the intrinsic (redshift corrected) peak energy ($E_{p,i}$) and the isotropic-equivalent gamma-ray energy release ($E_{iso}$). Lines show the best-fit correlation (solid) and their $3\sigma$ bounds (dashed) for the long (blue) and short (grey) GRB populations. Correlation fits and comparison data are from Minaev & Pozanenko (2020).
@@
Given that the GRBs in our sample were selected based on the recovery of their optical afterglows by a relatively shallow telescope like GOTO, it’s perhaps not surprising that they’re outliers with respect to the wider sample of GRBs. GRBs 240916A and 241228B both appear to be straightforward cases of intrinsically powerful GRBs, as evidenced by their higher than average $E_{iso}$. However, their measured $E_{p,i}$ is high relative to the Amati relation even when accounting for the large $E_{iso}$.

### v2

@@ 4 Afterglow Follow-up Observations
Figures 5(a–g) illustrate the timeline of the multi-wavelength follow-up campaigns for all GRBs in the sample, marking the epochs at which each facility recorded its first observation relative to the trigger time. The various phases of the events are colour-coded as follows: X-ray afterglow (blue), UV/optical/NIR afterglow (green), radio afterglow (purple), and prompt emission (red) for completeness. This timeline highlights the wide temporal coverage and the rapid, coordinated response from both ground- and space-based observatories across the electromagnetic spectrum for the GRBs in our sample. More observation details are provided below. A full log of photometric measurements, combining GCN Circular reports and data from this work, is presented in Tables LABEL:tab:phot_data and A2, while Table 3 summarises the spectroscopic campaigns.
@@ 4.2.1 Swift/UVOT
In addition to XRT, Swift simultaneously observes with its UV-Optical Telescope (UVOT; Roming et al. 2005). We obtained the resulting data from the UK Swift Science Data Centre [^6] https://www.swift.ac.uk/index.php (UKSSDC) and used uvotproduct v2.9777As part of HEASOFT v6.32 (Nasa High Energy Astrophysics Science Archive Research Center (Heasarc), 2014). to measure the photometry of the afterglow. We used a 5″ radius circular aperture centred at the positions noted in Section 3.2 and a detection threshold of $3\sigma$. The measured magnitudes were converted from the UVOT photometric system to AB using the standard UVOT zeropoints (Breeveld et al., 2011). The afterglow was detected in at least one epoch for five of the seven sources in our sample. The exceptions are GRB 240122A, where the UVOT FoV did not cover the source position, and GRB 240225B, for which no UVOT data were obtained. The UVOT follow-up observations are listed in Table LABEL:tab:phot_data, and the corresponding light curves are presented in Figure 6.
@@
The IO:O Imager at the robotic 2m Liverpool Telescope (LT; Steele et al., 2004) located at the international Observatorio del Roque de los Muchachos, La Palma, was triggered for follow-up of GRBs 240122A, 240225B, and 240910A. For GRBs 240122A and 240910A, one epoch was obtained in each of rizriz bands. For GRB 240225B, one epoch is in grizgriz, and three later epochs were secured in $r$-band. All LT data were pre-reduced for bias, dark, and flat-field corrections using the facility pipeline. The photometry was then extracted with the photometry-sans-frustration pipeline (psf; Nicholl et al., 2023), making use of its built-in template subtraction. The observations are summarised in Table LABEL:tab:phot_data, and the corresponding light curves are shown in Figure 6.
@@ 4.2.3 NOT
4.2.4 VLT/X-shooter acquisition camera
@@ 4.2.6 OSN
We processed all our observations using the APEX pipelines (Devyatkin et al., 2010; Kouprianov, 2012; Pankov et al., 2022). The process involved image calibration (dark frame subtraction, flat-fielding, cosmic-ray removal), image quality control, image stacking, and source extraction. The apex_forced_phot pipeline was utilised for the forced photometry of the GRB 240225B afterglow on difference images. Image subtraction with the method described in Tomaney & Crotts 1996, was performed by apex_subtract pipeline, using Pan-STARRS-DR1 survey images as a reference obtained from HIPS2FITS service [^11] https://alasky.cds.unistra.fr/hips-image-services/hips2fits (Boch et al., 2020). This step ensured that the underlying host galaxy, presented in the Legacy Survey DR9 with a magnitude of $r\sim 24.2$ and a photometric redshift of $z\sim 0.9$, did not affect the afterglow measurement.
@@ 4.2.11 Altay
The 1.6m Multi-channel Photometric Survey Telescope (Mephisto) is a wide-field multichannel telescope (Yuan et al., 2020). It is located at Lijiang Observatory of Yunnan Astronomical Observatories, Chinese Academy of Sciences, and is operated by the South-Western Institute for Astronomy Research, Yunnan University. Equipped with three-channel CCD cameras (blue $uv$, yellow $gr$, and red $iz$ channels), Mephisto can perform simultaneous observations in $ugi_{m}$ or $vrz_{m}$ optical bands at a particular moment. The wavelength coverage of the $u_{m},v_{m},g_{m},r_{m},i_{m}$, and $z_{m}$ filters is 320–365, 365–405, 480–580, 580–680, 775–900, and 900–1050 nm with central wavelengths at 345, 385, 529, 628, 835, and 944 nm, respectively (see e.g., Yang et al., 2024; Cheng et al., 2025). Presently, the facility is in an advanced stage of commissioning.
@@ 4.3 Spectroscopic Observations
For the GRBs in our sample, we acquired spectra using the X-shooter instrument mounted on the VLT (Vernet et al., 2011), and OSIRIS (Optical System for Imaging and low-Intermediate-Resolution Integrated Spectroscopy) on the Gran Telescopio Canarias (GTC; Cepa 1998), except for GRB 241002B, for which we did not get any spectroscopic observation due to scheduling constraints. A complete summary of the VLT/X-shooter and GTC/OSIRIS spectroscopic configurations, exposure details, and observing conditions is provided in Table 3, whereas spectra are shown in Figures A1A7.
@@ 4.4.1 ATCA
Following its optical localisation, GRB 240122A was observed with the ATCA under the PanRadio GRB programme C3542 (PI: Anderson) on 2024 January 24, 26, 28, and February 12. This program aims to perform high-cadence multi-frequency radio monitoring of a large sample of LGRBs in the southern hemisphere (Declinations <10<-10 deg) between minutes to years after the burst to explore the evolution and properties of their afterglows (Leung et al., in prep.; Anderson et al., in prep.). GRB 240122A was observed with a wide range of frequencies centred on 5.5, 9.0, 16.7, and 21.2 GHz, each with a 2048 MHz-wide band. We reduced the visibility data using standard routines in MIRIAD (Sault et al., 1995). We used a combination of manual and automatic RFI flagging before calibration, conducted with MIRIAD tasks uvflag and pgflag, respectively. We used PKS 1934-63 to determine the bandpass response and to calibrate the flux density scale for all frequency bands. We used PKS 0607-157 to calibrate the time-variable complex gains for all epochs and frequency bands. After calibration, we inverted the visibilities using a robust weighting of 0.5 and then used the CLEAN algorithm (Clark, 1980) to the target source field using standard MIRIAD tasks INVERT, CLEAN, and RESTOR to obtain the final images. For each observation, we measure the flux density of a detected source by fitting a point-source model to the restored image using the Miriad task imfit and report a non-detection using the rms sensitivity obtained from the residual image. The $1\sigma$ errors reported are purely statistical, as the systematic errors are expected to be much smaller ($\lesssim 5$%; e.g., Reynolds, 1994; Tingay et al., 2003). We detected the radio counterpart at 9 GHz on both 2024 January 24 and 28 at a position consistent with the GOTO optical counterpart (Kumar et al., 2024b). For all other frequencies, we estimated $3\sigma$ upper limits (see Table A2).
@@ 4.4.2 VLA
[…]
@@
GRB $z$ $T_{\rm 90}$ Model Fluence (2 – 20 keV) $\alpha$ $\beta$ $E_{\rm p}$ PG-Statistic DoF HR
@@
240122A $3.163$ $\approx 36$ PL $1.26^{+0.47}_{-0.94}$ $1.9\pm 0.4$ $1.32^{+1.30}_{-0.65}$
240225B $0.946$ $\approx 21$ PL $10.7^{+0.63}_{-1.26}$ $1.9\pm 0.1$ $1.32^{+0.25}_{-0.21}$
GRB $z$ $T_{\rm 90}$ Model Fluence (10 – 1000 keV) $\alpha$ $\beta$ $E_{\rm p}$ PG-Statistic DoF HR
@@ 5.1 Prompt Emission Analysis
Here, we describe the methods used to extract and analyse the prompt emission properties for each GRB, using data from the relevant high-energy instruments. The analysis is divided into two parts: MAXI/GSC events and the Fermi/GBM events.
@@ 5.1.1 GRBs 240122A and 240225B with MAXI/GSC
We analysed the prompt emission of GRBs 240122A and 240225B with archival data of MAXI/GSC using High Energy Astrophysics Software (HEASOFT [^14] https://heasarc.gsfc.nasa.gov/docs/software/heasoft/). X-ray events of gsc_med type are processed with mxproduct. Because light curves and spectra produced by mxproduct are not suitable for short and variable transients like GRBs, we performed an additional step to extract light curves with a 1-second time resolution and applied effective area correction. Note that the process is identical to the process used for the MAXI GRB catalogue.
@@
Then, we extracted a spectrum of the $T_{\rm 90}$ interval and corrected for variations in the effective area. The spectra of GRBs 240122A and 240225B are fit with a single power-law model, and the photon indices are found to be $1.9\pm 0.4$ and $1.9\pm 0.1$ ($1\sigma$ error), respectively. The energy flux in 2–20 keV was $3.5_{-2.6}^{+1.3}\times$10-9 erg cm-2 s-1 and $5.1_{-0.6}^{+0.3}\times$10-8 erg cm-2 s-1. Figure A9 shows the spectrum with the best-fitting model. The results of each fit are shown in Table 4.
@@ 5.1.2 GRBs 240619A, 240910A, 240916A, 241002B, and 241228B with Fermi/GBM
We analysed the prompt emission of GRBs 240619A, 240910A, 240916A, 241002B, and 241228B using the Fermi/GBM data available from the High Energy Astrophysics Science Archive Research Center (HEASARC [^15] https://heasarc.gsfc.nasa.gov) archive (Gruber et al., 2014; von Kienlin et al., 2014; Bhat et al., 2016; von Kienlin et al., 2020). Using the HEASOFT and the Fermi Gamma-ray Data Tools (Goldstein et al., 2023), we took the time-tagged event (TTE) data in the $T_{\rm 90}$ interval for each burst to use as the source. We used the data from the brightest NaI detectors and the corresponding BGO detectors. These were NaI 0, NaI 1, NaI 2 and BGO 0 for GRB 240619A; NaI 8, NaI 11, and BGO 1 for GRB 240910A; NaI 3, NaI 4, and BGO 0 for GRB 240916A; NaI 10, NaI 11, and BGO 1 for GRB 241002B; and NaI 6, NaI 7, and BGO 1 for GRB 241228B. The background was modelled in the standard way, using a polynomial function fit to the CSPEC data for each burst. Polynomial order increased until the reduced fit statistic was <1.15<1.15, resulting in the models fitting each GRB suitably well. Once the background model was obtained, we interpolated it across the source interval and exported this as the background to be used for spectral analysis. Additionally, we extracted the necessary response files.
@@ 5.2 Prompt Emission Properties
We plot the four of five Fermi bursts where it was possible to measure both $E_{\rm p}$ and $E_{iso}$ (see Table 4) on the Amati plane (Amati et al., 2002; Amati, 2006) in Figure 7. GRBs 240916A and 241228B appear to show unusually high redshift-corrected peak energies ($E_{p,i}$) relative to their $E_{iso}$ measurements, and are inconsistent with the Amati relation at the $3\sigma$ level. Both bursts lie at the high end of the $E_{iso}$ distribution, indicating intrinsically powerful GRBs. In addition to this, the high energy photon index ($\beta$) measured for GRB 240619A is less than 2, indicating that the true peak of the spectrum is at an even higher energy. This may make 240619A an outlier of the Amati relation too. Only GRB 240910A appears to be typical in terms of its measured prompt properties.
We also plot all five Fermi and two MAXI GRBs on an HR – $T_{\rm 90}$ diagram (cf. Kouveliotou et al., 1993), see Figure 8. It is immediately apparent that the Fermi GRBs are unusually spectrally hard. Only GRB 241002B (best fit with a power-law model with a photon index of $\alpha=1.97^{+0.08}_{-0.06}$) sits within the main ‘cloud’ of LGRBs. The others are either longer in $T_{\rm 90}$ (GRB 240910A), or driven to high HR by their abnormally hard (for collapsar GRBs) photon indices (GRBs 240619A, 241228B) and/or high $E_{\rm p}$ (GRB 240916A). The MAXI GRBs also sit at the high end of the GBM HR distribution, but these values are obtained by extrapolating the 2–20 keV spectral fits to 300 keV, and so should be considered upper limits because we have no constraints on any breaks in the spectrum.
Figure 7: Fermi GRBs in our sample plotted in the Amati plane (Amati et al., 2002; Amati, 2006), showing the relationship between the intrinsic (redshift corrected) peak energy ($E_{p,i}$) and the isotropic-equivalent gamma-ray energy release ($E_{iso}$). Lines show the best-fitting correlation (solid) and their $3\sigma$ bounds (dashed) for the long (blue) and short (grey) GRB populations. Correlation fits and comparison data are from Minaev & Pozanenko (2020).
@@
Given that the GRBs in our sample were selected based on the recovery of their optical afterglows by a relatively shallow telescope like GOTO, it is perhaps not surprising that they’re outliers with respect to the wider sample of GRBs. GRBs 240916A and 241228B both appear to be straightforward cases of intrinsically powerful GRBs, as evidenced by their higher than average $E_{iso}$. However, their measured $E_{p,i}$ is high relative to the Amati relation even when accounting for the large $E_{iso}$.